### v2

# RDB2G-Bench: A Comprehensive Benchmark for Automatic Graph Modeling of Relational Databases

Abstract
[…]
1 Introduction
A relational database (RDB) is a collection of data organized into multiple tables connected by shared keys. RDBs enable systematic and efficient management of related information through query languages such as SQL, and have been widely adopted across diverse industries, including finance [18], healthcare [34], and e-commerce [30]. This widespread use has led to the emergence of diverse machine learning applications built upon RDBs.
For machine learning on RDBs, recent studies have explored graph-based approaches [11, 12, 29, 35, 42, 6], which involve modeling RDBs as graphs. Typically, rows of tables are modeled as nodes in the resulting graph, and foreign key (FK) relationships are represented as edges. Graph modeling, where graph neural networks are subsequently applied, effectively captures structural dependencies and leads to improved performance across a range of machine learning tasks [8, 29, 35].
One important consideration in graph-based approaches is the variety of ways to model RDBs as graphs. For example, a single table row representing a user transaction on an item can be modeled either as (1) a node representing the transaction or (2) an edge linking the corresponding user and item. Additionally, one may choose to include only a subset of tables (i.e., nodes) and FK relationships (i.e., edges), excluding those considered less relevant. We call the process of representing RDBs as graphs RDB-to-graph modeling, and refer to the resulting graph representation as a graph model.
Careful RDB-to-graph modeling is crucial since the empirical effectiveness of graph-based approaches heavily depends on the choice of graph model. For example, our analysis of a graph-based approach [11, 29] on the rel-avito dataset shows that modeling the rows of certain tables as edges (rather than nodes) and excluding some FK relations led to a downstream-task performance gain of up to 5% compared to using a commonly-used fixed modeling rule [11], as illustrated in Figure 1.
Despite the importance of intelligent RDB-to-graph modeling, research in this area remains in its early stages, with one main reason being the difficulty of evaluation. First, there are numerous possible graph models (see Figure 2(a) for examples from real-world RDBs), making exhaustive comparisons extremely expensive. Moreover, evaluating even a single graph model is computationally costly, as it typically involves training graph neural networks (GNNs) on a large graph.
To facilitate research on intelligent RDB-to-graph modeling, we introduce RDB2G-Bench, the first RDB-to-graph modeling benchmark framework, with the following key contributions:
  • Datasets. RDB2G-Bench provides precomputed performance metrics (training, validation, and test performance and runtime over at least five trials) for 50k graph models based on $5$ real-world RDBs and 12 predictive tasks. They enable researchers to evaluate their modeling methods without training GNNs on the resulting graph models.
[…]
  • Observations. Our analysis of the datasets and benchmark results identifies important factors that impact performance, providing practical insights for effective graph modeling.
Figure 1: Overview of key concepts. An RDB schema is converted into various network schemas using different RDB-to-graph (RDB2Graph) modeling methods. Graphs are then constructed from these schemas, where graph neural networks (GNNs) are trained and evaluated. In the given example task, optimal modeling yields up to a 5% performance improvement over a widely-used heuristic [11]. Note that the optimal graph model selectively uses tables and foreign key (FK) relations, with table rows modeled as edges, while the heuristic models the entire RDB with all table rows as nodes.
Our datasets and code are available at: https://github.com/chlehdwon/RDB2G-Bench.
2 Related Works
2.1 Relational Deep Learning
Early approaches to relational database (RDB) learning relied on feature engineering [19] (e.g., table joins), transforming an RDB into a single table to leverage tabular machine learning [5, 21, 2, 13]. Recent works on relational deep learning (i.e., deep learning on RDBs) [8, 31, 11, 12, 9] pioneered modeling RDBs as graphs, which facilitate the capture of cross-table dependencies. Building on these graph models, advanced methods [6, 42] have achieved improved predictive performance on RDBs.
2.2 Benchmarks on Relational Deep Learning
Early benchmarks (e.g., the CTU relational learning repository [26]) set the stage for evaluating machine learning (ML) models on real-world RDBs. Recent benchmarks (e.g., RDBench [43], RelBench [29], and 4DBInfer [35]) shift the focus to graph-based approaches, which often show improved performance over traditional tabular learning methods [21, 5].
Distinctive property of our benchmark: Existing benchmarks provide RDBs and predictive tasks to compare machine learning (ML) methods, which are typically based on the same graph presentation of the RDBs. Our benchmark, RDB2G-Bench, however, is designed to evaluate RDB-to-graph modeling strategies, which provide graph models based on which various modeling methods perform.
Specifically, by offering precomputed performance metrics for 50k graph models, RDB2G-Bench enables comparing various graph modeling strategies without repeated GNN training (see Section 5.2 for efficiency gains).
2.3 Graph-based Modeling of RDBs
Traditionally, ML on RDBs has relied heavily on heuristic methods for automating schema transformations. These methods include rule-based relationship mining [1, 40, 25, 23] and conversion of RDBs into single-table formats through predefined aggregation functions [20, 44]. In addition, recent studies have focused on graph modeling, such as mapping table rows to nodes (Row2Node) [29] or to nodes or edges (Row2N/E) [35]. However, heuristic-driven approaches often produce suboptimal RDB representations (see Figure 1 for an example). More recently, AutoG [7] was proposed, leveraging large language models (LLMs) to actively explore effective graph models of RDBs.
Our benchmark, RDB2G-Bench, enables the evaluation of various RDB-to-graph modeling methods, including both heuristic and LLM-based approaches (see Section 5.1 for the ten methods included in the benchmark), ultimately fostering the development of more advanced modeling strategies.
3 Dataset Design for RDB2G-Bench
In this section, we introduce the datasets provided in RDB2G-Bench, our benchmark for the RDB-to-graph modeling problem. The problem is defined as follows:
Problem Definition.
Definition 1 (RDB-to-Graph Modeling).
Given: A relational database $\mathcal{R}$, and a graph neural network $\mathcal{M}$ for a downstream task $\mathcal{T}$. To Find: the graph model $\mathcal{G}^{*}$ of $\mathcal{R}$ that maximizes the performance $\mathcal{P}$ of $\mathcal{M}$ trained and evaluated on $\mathcal{T}$, i.e., $\mathcal{G}^{*}=\arg\max\nolimits_{\mathcal{G}\in\mathcal{F}(\mathcal{R},\mathcal{T})}\mathcal{P}(\mathcal{M}(\mathcal{G}),\mathcal{T}),$ where $\mathcal{F}(\mathcal{R},\mathcal{T})$ is the set of possible graph models of $\mathcal{R}$ on $\mathcal{T}$, defined in Section 3.1; and example metrics of $\mathcal{P}$ are provided in Section 3.2.
Each dataset comprises graph models (i.e., graph-structured representations) of a given relational database (RDB) paired with downstream task performance metrics of graph neural networks (GNNs) (denoted by $\mathcal{M}$) trained on those graph models. We first describe the considered design space of graph models (Section 3.1), followed by the procedure for obtaining GNN performance metrics on each graph model (Section 3.2)
RDBs and downstream predictive tasks. Our RDB2G-Bench datasets were created based on 5 real-world RDBs and 12 predictive tasks provided in RelBench [29]. The RDBs and tasks, summarized in Figure 2(a), were carefully selected to cover a diverse set of tasks (classification, regression, and recommendation) and domains. Refer to Appendix A.1 for the details of the RDBs and tasks.
3.1 Construction of Graph Models
As discussed in Section 1, there are various ways to model each RDB as a graph, and the performance of GNNs depends heavily on the chosen graph model. In RDB2G-Bench, we consider the design choices in the following two steps of RDB-to-graph modeling, both of which have a significant impact on the downstream-task performance of GNNs, as shown in Section 4:
  • Step 1: Selecting which tables and foreign key (FK) relationships to include in the graph model.
  • Step 2: Selecting how to represent the rows of each table, as either nodes or edges, in the graph.
That is, for each RDB $\mathcal{R}$ (e.g., rel-avito), our dataset consists of graph models that result from different combinations of design choices in Step 1 and Step 2, which we denote by $\mathcal{F}(\mathcal{R},\mathcal{T})$ in Definition 1 (refer to Figure 1 for examples). These graph models are used to train graph neural networks (GNNs) for predictive tasks (e.g., user-clicks) defined on the RDB (as described in Section 3.2), and to be used for this purpose, the graph models must satisfy several constraints. First, a valid graph model must select the task table, which the considered predictive task is defined on. Second, all selected tables should be connected to the task table via a path whose length does not exceed the number of GNN layers. Violating this constraint would result in some nodes being unreachable during message passing, leading to degraded performance. Third, (the rows of) a table can be modeled as edges only if the table has exactly two FKs and its primary key (PK) is not referenced by any FKs in other tables. Note that, for a table with more than two FKs, hyperedge modeling would be required, which is beyond our scope, and edge modeling of a table whose PK is referenced by FK renders those FK relationships unrepresentable. Also note that these constraints may lead to different graph model spaces for downstream tasks defined on the same RDB. As summarized in Figure 2(a), we constructed about 50k graph models spanning the aforementioned RDBs and downstream tasks; and evaluated their downstream task performance metrics, as described below.
(a) Summary of the RDB2G-Bench datasets, which cover 50k graph models in total.
RDB Task Name Type # Tables # Graph Performance Statistics
Models Best AR2N [29] Worst
rel-avito user-clicks (UC) classification 8 944 67.93 64.66 60.89
user-visits (UV) classification 944 66.33 65.97 59.83
ad-ctr (AC) regression 1304 0.039 0.040 0.044
user-ad-visit (UAV) recommendation 909 3.682 3.661 0.159
rel-event user-repeat (UR) classification 5 214 82.29 77.65 63.96
user-ignore (UI) classification 214 82.82 82.22 74.29
user-attendance (UA) regression 214 0.237 0.244 0.266
rel-f1 driver-dnf (DD) classification 9 722 74.56 73.14 67.40
driver-top3 (DT) classification 722 81.88 78.11 75.37
driver-position (DP) regression 722 3.831 3.913 4.171
rel-stack post-post-related (PPL) recommendation 7 7979 12.04 10.82 0.006
rel-trial study-outcome (SO) classification 15 36863 70.91 68.09 62.85
(b) Visualization of the RDB2G-Bench datasets on three predictive tasks.
Figure 2: (a) We summarize the RDBs, tasks, and their associated graph models. For each classification, regression, and recommendation task, we collect AUC-ROC (%), MAE, and MAP (%), respectively, on each graph model. For each task, we report the performances on the best graph model, the worst model, and that given by AR2N modeling [29]. (b) For three tasks (driver-top3, user-attendance, post-post-related), we visualize the distribution of performances on the downstream task (Y-axis) across all graph models, along with training time per epoch (X-axis) and the parameter size of the graph neural network (indicated by color). Note that there exist graph models yielding substantial improvements in both performance and efficiency compared to those generated by widely-used AR2N modeling [29].
3.2 Collection of Performance Metrics
For each task and each constructed graph model, we collected the following performance metrics (i.e., $\mathcal{P}$ in Definition 1):
  • Predictive performance: The training, validation, and test performances on the downstream task.
  • Runtime: The total elapsed time per epoch over at least five trials.
  • Parameter size: The total number of learnable parameters of predictive GNNs, which depends on the graph model.
Specifically, we collected the performance metrics under the following setups:
Machines. All experiments were conducted using NVIDIA RTX A6000 GPUs with 48GB of memory and Intel Xeon Silver 4210/4310 CPUs. Constructing all datasets required about 10,400 GPU hours.
Predictive GNNs. As predictive machine learning models, we used GNNs provided by RelBench [29], which leverage PyTorch-Frame [17] to encode each table as input to the GNNs. Specifically, for classification and regression tasks, we employed Heterogeneous GraphSAGE [14] with sum aggregation to update embeddings for the final predictions. For recommendation tasks, we utilized ID-GNN [41]. As the RelBench GNNs do not support edge modeling of tables, we extended them to incorporate edge features and used them when tables are modeled as edges (refer to Appendix A.2). We additionally used three alternative predictive GNNs, as described in Section 4.5.
Training Details. Based on the training protocol provided by RelBench [29] (including the training, validation, and test splits), we tuned only the learning rate for each graph model. The other hyperparameters were fixed to the combinations that yielded the best overall performance on 50 randomly sampled graph models. For classification and regression tasks, we selected the learning rate from {0.005, 0.001, 0.0005}, and for recommendation tasks, from {0.001, 0.0005, 0.0001}. The number of training epochs was fixed at 20, which was empirically sufficient for convergence. For reliability, we repeated each experiment with 15 different random seeds for the rel-f1 and rel-event datasets, and with 5 seeds for the remaining datasets. Refer to Appendix A.3 for further details.
4 Observations from the RDB2G-Bench Datasets
In this section, we highlight key observations from our analysis of the constructed RDB2G-Bench datasets. Unless otherwise stated, we use the predictive models described in Section 3.2. For the full analysis results omitted due to space constraints, refer to Appendix B.
4.1 Obs 1. “Finding the best graph models is worthwhile, given their performance benefits.”
Analysis Overview. We analyze the distribution of downstream-task performances across graph neural networks (GNNs) trained on different graph models of RDBs. We also examine how performance relates to (1) mean GNN training time (per epoch) and (2) the number of GNN parameters.
Results. As shown in Figure 2(b), the performance varies significantly across graph models. For example, in the post-post-related task, the performance difference spans 11.3%. Notably, for each task, there exist graph models that lead to significantly better performance (e.g., 4.83% improvement in the driver-top3 task) than the one produced by the widely used all-rows-to-nodes (AR2N) modeling [29], and they often even require shorter training time and fewer parameters.
Implications. These results suggest that it is not always beneficial to include all tables and foreign key (FK) relationships in a graph model. Instead, selecting only the most relevant tables can lead to better performance, while also reducing GNN parameters and training time. Moreover, there exist graph models that are both effective and efficient, yet difficult to identify using simple heuristics.
4.2 Obs 2. “Modeling table rows as edges can be crucial, depending on the task.”
Analysis Overview. We compare the two design choices for modeling the rows of a table in an RDB: (1) Row2Node, which always represents table rows as nodes, and (2) Row2Edge, which represents table rows as edges, especially when the table satisfies the constraints discussed in Section 3.1. Specifically, we compare the downstream task performance distributions of the top 10% highest-performing graph models, grouped by design choice.
Results. Figure 3 highlights two representative patterns. For the user-ignore task on the rel-event dataset, Row2Edge modeling consistently leads to performance degradation across all tables. In contrast, for the user-repeat task on the same RDB, the Row2Edge modeling of the EA (event_attendees) and UF (user_friends) tables significantly improves AUC-ROC scores.
Implications. These findings suggest that even within the same RDB, the benefit of Row2Edge modeling varies significantly across downstream tasks. This variability suggests that no universal rule-of-thumb exists for modeling table rows; the choice should be carefully designed to achieve good performance on the target downstream tasks.
4.3 Obs 3. “Top-performing graph models share common substructures.”
Analysis Overview. We investigate whether top-performing graph models share common substructures by examining the top 1% best-performing graph models for each task.
Results. Figure 4 shows a representative case from the user-attendance task on the rel-event dataset. Note that all top-5 best performing graph models include the foreign-key relationship events $\rightarrow$ users. They also model either the event_attendees table or the event_interest table as edges.
Implications. This empirical observation indicates the presence of common substructures that are critical for downstream task performance. Finding these structures can be key to identifying effective and efficient graph models.
Figure 3: Modeling table rows as edges (Row2Edge) can be crucial, depending on the task (Obs 2). EA (event_attendees), EI (event_interest), and UF (user_friends) indicate the tables whose rows can be modeled as edges. Note that Row2Edge modeling improves performance for the user-repeat task, but not for the user-ignore task, even when both are defined on the same RDB.
Figure 4: Top-performing graph models share common substructures (Obs 3). As shown in their graph models, the top-5 graph models commonly (a) include the foreign-key (FK) relationship events $\rightarrow$ users and (b) model either the event_attendees table or the event_interest table as edges. Note that the users_friends table has two FKs (user and friend) both referencing the users table.
4.4 Obs 4. “Different tasks may require different graph models, even on the same RDB.”
Analysis Overview. We analyze cross-task performance correlations to examine whether the effectiveness of a graph model in one downstream task implies its effectiveness in others. Specifically, we measure the Spearman rank correlation [33] between performances on tasks within the same RDB.
Results. As shown in Figure 5, tasks whose objectives are highly aligned (spec., user-attendance and user-repeat in rel-event)111user-attendance asks how many events each user will attend, and user-repeat asks whether a user will repeat attendance at an event. exhibit a high Spearman rank correlation exceeding 0.9. However, most task pairs exhibit low correlations (below 0.4), indicating that a graph model effective for one task may not generalize well to others, even within the same RDB.
Implications. These results suggest that an effective RDB-to-graph modeling strategy should account not only for the characteristics of the RDB but also for those of the specific downstream task. Moreover, reusing a graph model is effective only across tasks with closely aligned objectives.
Figure 5: Different tasks may require different graph models, even on the same RDB (Obs 4). Spearman correlations between downstream task performances on each RDB (rel-f1, rel-event, or rel-avito) are generally low (below 0.4), except for tasks with closely aligned goals.
4.5 Obs 5. “Effectiveness of graph models generalizes across predictive GNNs.”
Analysis Overview. We analyze performance correlations across different predictive GNNs applied to the same graph models to examine whether a graph model effective with one GNN (e.g., Heterogeneous GraphSAGE [14]) is also effective with others. To this end, we consider three additional GNNs (GraphSAGE with mean aggregation [14] and GIN [39]) and a graph transformer (GPS [27]) For their details, refer to Appendix A.2. As in the previous analysis, we employ the Spearman rank correlation [33].
Results. As shown in Figure 6, the correlations are generally high (above 0.7) in most cases, indicating that the effectiveness of graph models tends to generalize across different GNNs. Notably, the graph transformer also exhibits strong correlations with GNNs, despite their architectural differences.
Implications. This cross–GNN generalizability of graph models suggests that RDB-to-graph modeling strategies trained and shown to work well using our benchmark, RDB2G-Bench, can be effective across various predictive GNNs, highlighting the broad utility of RDB2G-Bench.
Figure 6: Effectiveness of graph models generalizes across predictive GNNs (Obs 5). Spearman correlations between different predictive GNNs—GraphSAGE with sum and mean aggregation (denoted as sum and mean), GIN, and GPS—are generally high (above 0.8) across the three RDBs (rel-avito, rel-trial, and rel-event).
5 Benchmark Results on the RDB2G-Bench Datasets
In this section, we review the benchmark results of ten RDB-to-graph modeling methods on our RDB2G-Bench datasets.
5.1 Baselines for RDB-to-Graph Modeling
We benchmark ten RDB-to-graph modeling methods, which are categorized into (a) heuristic-based methods, (b) action-driven search algorithms, and (c) an LLM-based approach. All methods are designed to select a graph model among those in our RDB2G-Bench datasets (see Section 3.1 for details on our graph model space).
Heuristic-Based Methods. The following heuristic-based methods rely on simple rules to model an RDB as a graph, without explicitly searching for effective graph models:
  • S1. Random [4, 24]: It randomly samples graph models up to the budget and selects the one with the highest downstream-task performance.
  • S2. All-Rows-to-Nodes (AR2N) [29]: This widely-used method includes all tables and foreign-key (FK) relationships in the graph, with all table rows modeled as nodes, as shown in Figure 1.
Action-Based Search Algorithms. Action-based search algorithms explore and optimize graph models by iteratively modifying them using a predefined set of operations, referred to as actions.
We consider the following four actions, which are designed to effectively span our graph model space: (1) add_fk_edge: adding an FK relationship between tables to the graph model, (2) remove_fk_edge: removing an existing FK relationship, (3) convert_row_to_edge: changing the modeling of a table from Row2Node to Row2Edge, and (4) convert_edge_to_row: changing the modeling of a table from Row2Edge to Row2Node. We build the following six methods based on these actions:
  • S3. Greedy Forward Search (GF): It starts from the target table and greedily repeats an action (except for remove_fk_edge) that yields the greatest improvement to the current graph in terms of downstream-task performance. Recall that tables that are disconnected or distant from the target table are excluded from the graph model, as described in Section 3.1.
  • S4. Greedy Backward Search (GB): It starts from the AR2N graph, which includes the entire RDB, and greedily repeats an action (except for add_fk_edge) that yields the greatest improvement to the current graph.
  • S5. Greedy Local Search (GL): It starts from a random graph and greedily repeats an action of any type that provides the greatest improvement to the current graph.
  • S6. Evolutionary Algorithm (EA): It applies evolutionary principles, including mutation and selection, to iteratively evolve graph models over generations. It randomly applies the predefined actions of any type to mutate the current graph and search for improved ones. Our implementation follows the regularized evolution strategy [28].
  • S7. Bayesian Optimization (BO): It applies a Bayesian optimization algorithm, specifically BANANAS [37], to the RDB-to-graph modeling task. This approach efficiently explores the graph model space by (a) modeling the function that maps actions to performance and (b) iteratively selecting actions of any type that are likely to improve the graph.
  • S8. Reinforcement Learning (RL): It applies a controller based on recurrent neural networks [32, 15] and trains it via a policy gradient descent approach [38, 3]. The RL-based predictor learns to select actions of any type at each step of the RDB-to-graph modeling process to optimize downstream-task performance.
Further details of the action-based search algorithms are provided in Appendix C.
LLM-Based Baselines. Inspired by AutoG [7], we developed the following two LLM-based baselines that generate action sequences for RDB-to-graph modeling using LLMs.
  • S9. LLM baseline (LLM): It directly generates action sequences that specify how a given RDB is expressed as a graph based on the given prompts. At each step, an LLM is provided with a set of candidate actions and selects a sequence of actions expected to be most effective, based on its reasoning capabilities. The key distinction from AutoG lies in the simplified prompt design. Prompts irrelevant due to differences in action spaces are omitted.
  • S10. LLM-CoT baseline (LLM-CoT): It applies Chain-of-Thought [36] prompt design to encourage complex reasoning in the LLM baseline during the action selection step.
For fair comparisons with the above action-based search algorithms, both baselines operate over the same predefined set of action types. Claude Sonnet-3.5 is used as the backbone LLM, and details of the prompt designs are provided in Appendix F.
5.2 Benchmark Results
(a) Rank Averaged over Tasks
(b) Evaluation Cost
[…]
For fair comparisons, we evaluate all baselines under a budget, which limits the number of graph models whose effectiveness (e.g., ground-truth downstream-task performance) can be measured. We track the effectiveness of resulting graph models as the budget increases up to 5% of the total search space. Full results, omitted here due to space constraints, are provided in Appendix D.
Figure 7(a) shows the ranks of the baselines, averaged over 12 predictive tasks from RDB2G-Bench and 10 independent runs per baseline (3 runs for the LLM-based method). In addition, Table 1 shows the details results on the rel-f1 dataset. Based on these results, we derive several key observations:
  • Most baselines outperform AR2N modeling [29] with minimal exploration, suggesting that actively searching for effective graph models is preferable to relying on a fixed modeling rule.
  • Greedy methods and Random perform comparably to advanced approaches such as Bayesian Optimization and Reinforcement Learning. Especially, Greedy Backward performs best under small budgets, while Bayesian Optimization outperforms it as the budget increases. Therefore, the choice between them can be guided by the available computational resources or time constraints.
  • Value-based optimization algorithms, such as Bayesian Optimization and Greedy Forward, typically improve in effectiveness as performance feedback accumulates. However, under limited compute budgets, more complex algorithms, such as Reinforcement Learning, or algorithms affected by poor initialization, such as Greedy Local, often suffer from unstable exploration, resulting in poor performance.
  • In the short run, the LLM-based approach demonstrated strong capabilities in rapidly improving performance through contextual reasoning and effective multi-turn planning. However, it struggled to devise effective long-term action plans, resulting in limited gains as the budget increased. CoT reasoning shows slight performance improvements over the LLM baseline in most cases, which provides potential for using LLMs in RDB-to-graph transformation.
The strong performance of simple methods like greedy approaches indicates significant room for improving RDB-to-graph modeling.
[…]
Table 1: Performance of ten RDB-to-graph modeling methods on the rel-f1 dataset under varying budget levels. Refer to Appendix D.1 for results on other datasets.
Task Name driver-dnf (AUC-ROC (%) $\uparrow$) driver-top3 (AUC-ROC (%) $\uparrow$) driver-position (MAE $\downarrow$)
Methods Budget (%) Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 74.557 74.557 74.557 74.557 74.557 81.879 81.879 81.879 81.879 81.879 3.8311 3.8311 3.8311 3.8311 3.8311
Random 73.225 73.450 73.592 73.745 73.755 79.627 80.165 80.425 80.438 80.604 3.8498 3.8435 3.8420 3.8405 3.8399
AR2N 73.140 73.140 73.140 73.140 73.140 78.106 78.106 78.106 78.106 78.106 3.9125 3.9125 3.9125 3.9125 3.9125
GF 73.461 74.557 74.557 74.557 74.557 81.879 81.879 81.879 81.879 81.879 3.8591 3.8409 3.8409 3.8409 3.8409
GB 73.254 74.040 74.394 74.394 74.394 80.148 80.558 80.558 80.558 80.558 3.8437 3.8437 3.8437 3.8437 3.8437
GL 73.023 73.645 73.726 73.775 73.775 79.172 79.292 79.299 79.299 79.299 3.8649 3.8512 3.8500 3.8500 3.8500
EA 73.082 73.393 73.590 73.786 73.929 79.153 79.774 80.116 80.273 80.549 3.8470 3.8451 3.8423 3.8394 3.8394
BO 73.328 73.624 73.837 73.919 74.070 79.424 79.799 80.127 80.152 80.348 3.8520 3.8404 3.8399 3.8394 3.8388
RL 73.332 73.526 73.602 73.657 73.986 79.043 79.385 79.439 79.906 80.084 3.8542 3.8449 3.8417 3.8411 3.8411
LLM 73.857 73.857 73.857 73.857 73.857 80.337 80.536 80.536 80.607 80.607 3.8534 3.8534 3.8500 3.8500 3.8500
LLM-CoT 73.338 73.574 73.882 73.882 73.882 78.660 78.693 78.693 78.817 78.817 3.8437 3.8437 3.8437 3.8437 3.8437
6 Conclusions
[…]
Acknowledgements
This work was partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00406985, 20%). This work was partly supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00438638, EntireDB2AI: Foundations and Software for Comprehensive Deep Representation Learning and Prediction on Entire Relational Databases, 50%) (No. RS-2024-00457882, AI Research Hub Project, 10%) (RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale, 10%) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), 10%).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We reflect all the contributions and scope in the abstract and introduction(Section  1). In detail, we provide 3 contributions while describing the main claim of our paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We describe our limitations in Appendix E.1.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: Our paper doesn’t include any theoretical results.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Our results are reproducible because all experiment results are computed from RDB2G-Bench, which we provide in this paper. We also release our datasets and codes in public (Github, Huggingface).
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We provide all the codes and datasets through RDB2G-Bench. We also construct documents to faithfully reproduce the main experimental results. You can check all the code and data from the following URLs:
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We mentioned our hyperparameter settings in Section 3.2 and Section 5.1. Also all the details are described in Appendix A.3 and Appendix C.1.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We report standard deviations in Figure 1 and error bars in Figure 2, and leverage box plots in Figure 3 to justify the statistical significance of our experiments.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We describe all the computer resources and time of execution in Section 3.2 (See Machines.).
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We strictly follow the NeurIPS Code of Ethics, particularly the source of our dataset, which RelBench has already anonymized all the data. Also, we don’t contain any privacy information in our datasets and codes.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss various broader impacts of our paper in Appendix E.2.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We believe we do not have risks of misusing.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We used the datasets and models of RelBench [29] implementation, which mentioned several times in our paper. Also, there is no license violation in our paper.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We release RDB2G-Bench which includes datasets and codes. We provide the documentation to help users that can be found in our GitHub link: https://github.com/chlehdwon/RDB2G-Bench, and Hugging Face link: https://huggingface.co/datasets/kaistdata/RDB2G-Bench.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: We don’t use any crowdsourcing or human-related tasks in our paper.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: We do not utilize any crowdsourcing or human-related tasks.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [Yes]
  64. Justification: We implemented LLM-based method as a benchmark baseline, which motivated by AutoG [7] paper. The differences between AutoG are described in Section 5.1, and all the prompts that we used are in Appendix F.
Appendix A Dataset Detail of the RDB2G-Bench
Table 2: Statistics of RelBench.
RDB Task Name Type #Rows of Training Table #Unique Entities %Train/Test Entity Overlap
Train Validation Test
rel-avito user-clicks classification 59,454 21,183 47,996 66,449 45.3
user-visits classification 86,619 29,979 36,129 63,405 64.6
ad-ctr regression 5,100 1,766 1,816 4,997 59.8
user-ad-visit recommendation 86,616 29,979 36,129 63,402 64.6
rel-event user-repeat classification 3,842 268 246 1,514 11.5
user-ignore classification 19,239 4,185 4,010 9,799 21.1
user-attendance regression 19,261 2,014 2,006 9,694 14.6
rel-f1 driver-dnf classification 11,411 566 702 821 50.0
driver-top3 classification 1,353 588 726 134 50.0
driver-position regression 7,453 499 760 826 44.6
rel-stack post-post-related recommendation 5,855 226 258 5,924 8.5
rel-trial study-outcome classification 11,994 960 825 13,779 0.0
A.1 RelBench Detail
Our research involves five datasets and twelve tasks from RelBench [29], each representing diverse domains and varying scales. A detailed description of each dataset and task is provided below. For more details, please refer to the URLs. Additional statistics are presented in Table 2.
rel-avito  [^2] https://relbench.stanford.edu/datasets/rel-avito/.
Avito is a leading online advertisement platform, providing a marketplace for users to buy and sell a wide variety of products and services, including real estate, vehicles, jobs, and goods. The Avito Context Ad Clicks dataset on Kaggle is part of a competition aimed at predicting whether an ad will be clicked based on contextual information. This dataset includes user searches, ad attributes, and other related data to help build predictive GNNs.
  1. 1.
    user-visits: Predict whether each customer will visit more than one Ad in the next 4 days.
  2. 2.
    user-clicks: Predict whether each customer will click on more than one Ads in the next 4 days.
  3. 3.
    ad-ctr: Assuming the Ad will be clicked in the next 4 days, predict the Click-Through-Rate (CTR) for each Ad.
  4. 4.
    user-ad-visit: Predict the list of ads a user will visit in the next 4 days.
rel-event  [^3] https://relbench.stanford.edu/datasets/rel-event/.
The Event Recommendation database is obtained from user data on a mobile app called Hangtime. This app allows users to keep track of their friends’ social plans. The database contains data on user actions, event metadata, and demographic information, as well as users’ social relations, which captures how social relations can affect user behavior.
  1. 1.
    user-repeat: Predict whether a user will attend an event(by responding yes or maybe) in the next 7 days if they have already attended an event in the last 14 days.
  2. 2.
    user-ignore: Predict whether a user will ignore more than 2 event invitations in the next 7 days
  3. 3.
    user-attendance: Predict how many events each user will respond yes or maybe in the next seven days.
rel-f1  [^4] https://relbench.stanford.edu/datasets/rel-f1/.
The F1 database tracks all-time Formula 1 racing data and statistics since 1950. It provides detailed information for various stakeholders including drivers, constructors, engine manufacturers, and tyre manufacturers. Highlights include data on all circuits (e.g. geographical details), and full historical data from every season. This includes overall standings, race results, and more specific data like practice sessions, qualifying positions, sprints, and pit stops.
  1. 1.
    driver-dnf: For each driver predict the if they will DNF (did not finish) a race in the next 1 month.
  2. 2.
    driver-top3: For each driver predict if they will qualify in the top-3 for a race in the next 1 month.
  3. 3.
    driver-position: Predict the average finishing position of each driver all races in the next 2 months.
rel-stack  [^5] https://relbench.stanford.edu/datasets/rel-stack/.
Stack Exchange is a network of question-and-answer websites on topics in diverse fields, each site covering a specific topic, where questions, answers, and users are subject to a reputation award process. The reputation system allows the sites to be self-moderating.
  1. 1.
    post-post-related: Predict a list of existing posts that users will link a given post to in the next two years.
rel-trial  [^6] https://relbench.stanford.edu/datasets/rel-trial/.
The clinical trial database is curated from AACT initiative, which consolidates all protocol and results data from studies registered on ClinicalTrials.gov. It offers extensive information about clinical trials, including study designs, participant demographics, intervention details, and outcomes. It is an important resource for health research, policy making, and therapeutic development.
  1. 1.
    study-outcome: Predict if the trials will achieve its primary outcome.
A.2 GNN Implementation
In this section, we demonstrate the implementations of GraphSAGE, GIN, and GPS, including message passing and edge feature extensions for Row2Edge modeling. All implementations are based on the PyTorch Geometric (PyG) [10] library.
GraphSAGE Implementation. The standard GraphSAGE [14] message passing with sum aggregation is defined as:
$\mathbf{h}_{i}^{(k)}=\sigma\left(\mathbf{W}^{(k)}\cdot\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}^{(k-1)}+\mathbf{W}_{r}^{(k)}\cdot\mathbf{h}_{i}^{(k-1)}\right)$ (1)
Where, $\mathbf{h}_{i}^{(k)}$ is embedding of node $i$ at layer $k$, $\mathcal{N}(i)$ represents set of neighboring nodes of $i$, and $\sigma$ denotes non-linear activation function (e.g., ReLU).
To extend this formulation to incorporate edge features, $\mathbf{e}_{j\rightarrow i}$ for Row2Edge [35], we concatenate them with neighbor features. Then, Row2Edge message passing becomes as follows (the modified part is highlighted in red):
$\mathbf{h}_{i}^{(k)}=\sigma\left(\mathbf{W}^{(k)}\cdot\sum_{j\in\mathcal{N}(i)}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left[\mathbf{h}_{j}^{(k-1)}\,\|\,\mathbf{e}_{j\rightarrow i}\right]}+\mathbf{W}_{r}^{(k)}\cdot\mathbf{h}_{i}^{(k-1)}\right)$ (2)
GraphSAGE message passing with mean aggregation, which used in Section 4.5 is defined as:
$\mathbf{h}_{i}^{(k)}=\sigma\left(\mathbf{W}^{(k)}\cdot\frac{1}{N}\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}^{(k-1)}+\mathbf{W}_{r}^{(k)}\cdot\mathbf{h}_{i}^{(k-1)}\right)$ (3)
Where $N$ is defined as the number of neighbors.
GIN Implementation. Graph Isomorphism Network (GIN) [39] message passing is defined as:
$\mathbf{h}_{i}^{(k)}=\mathrm{MLP}^{(k)}\left((1+\epsilon^{(k)})\cdot\mathbf{h}_{i}^{(k-1)}+\sum_{j\in\mathcal{N}(i)}\mathbf{h}_{j}^{(k-1)}\right)$ (4)
Here, $\mathbf{h}_{i}^{(k)}$ denotes the embedding of node $i$ at layer $k$, $\epsilon^{(k)}$ is a trainable scalar parameter, and $\mathrm{MLP}^{(k)}$ is a Multi-Layer Perceptron (MLP) network used to transform aggregated messages. In our implementation, we adopt the following form:
$\mathrm{MLP}^{(k)}(\cdot)=\mathrm{Linear}(d,2d)\rightarrow\mathrm{ReLU}\rightarrow\mathrm{Linear}(2d,d)$ (5)
where $d$ is the hidden dimensionality of node features.
To incorporate edge features $\mathbf{e}_{j\rightarrow i}$, we adopt the GINE [16] formulation. Specifically, neighbor embeddings are modulated with their corresponding edge features before aggregation, as follows (the modified part is highlighted in red):
$\mathbf{h}_{i}^{(k)}=\mathrm{MLP}^{(k)}\left((1+\epsilon^{(k)})\cdot\mathbf{h}_{i}^{(k-1)}+\sum_{j\in\mathcal{N}(i)}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{ReLU}(\mathbf{h}_{j}^{(k-1)}+\mathbf{e}_{j\rightarrow i})}\right)$ (6)
GPS Implementation. The General, Powerful, Scalable Graph Transformer (GPS) [27] integrates local message passing with global attention, enabling expressive and scalable graph learning. GPS combines two components at each layer: (1) local neighborhood aggregation, and (2) global message exchange via Transformer-style self-attention.
Formally, for each layer $k$, the message passing is performed in two steps:
$\mathbf{h}_{i}^{(k)}=\mathrm{LocalMP}^{(k)}\left(\{\mathbf{h}_{j}^{(k-1)}:j\in\mathcal{N}(i)\}\right)+\mathrm{GlobalAttn}^{(k)}\left(\{\mathbf{h}_{j}^{(k-1)}:j\in\mathcal{V}\}\right)$ (7)
where $\mathbf{h}_{i}^{(k)}$ is the embedding of node $i$ at layer $k$, $\mathcal{N}(i)$ denotes the set of neighboring nodes of $i$, and $\mathcal{V}$ is the set of all nodes in the graph.
In our implementation, we employ Equation (1) and Equation (2) for the $\mathrm{LocalMP}^{(k)}$ of Row2Node and Row2Edge modeling, respectively, to naturally extend our settings to transformer-based architectures. To fit the GPS within a single GPU, we reduce the number of sampled neighbors from 128 to 32.
A.3 Hyperparameter Details
Table 3 shows our learning rate choices, which are only tuned for each task. For the other hyperparameter, all tasks used two-layer GNNs with a batch size of 512, 128 dimensions, and the Adam [22] optimizer.
Specifically, neighbor sampling counts for subgraph extraction are set to 128 when all tables are represented as nodes, and 12 when some tables are represented as edges. A lower sampling count is employed when tables are modeled as edges, since the two-layer GNN effectively captures paths longer than two hops, resulting in a significantly larger graph scale that cannot be processed by a single GPU.
Table 3: The learning rate settings tuned for each task.
RDB Task name Type Learning Rate
rel-avito user-clicks classification 0.001
user-visits classification 0.001
ad-ctr regression 0.0005
user-ad-visit recommendation 0.001
rel-event user-repeat classification 0.005
user-ignore classification 0.005
user-attendance regression 0.005
rel-f1 driver-dnf classification 0.005
driver-top3 classification 0.005
driver-position regression 0.0005
rel-stack post-post-related recommendation 0.0005
rel-trial study-outcome classification 0.0005
Appendix B Additional Results regarding Our Observations on the RDB2G-Bench Datasets
B.1 Additional Results regarding Obs 1.
Figure 8 complements Section 4.1 by presenting the performance distribution for the remaining 9 tasks. The results for driver-top3, user-attendance, and post-post-related are visualized in Figure 2 (b) in the main paper.
Figure 8: For the nine remaining tasks, we visualize the distribution of performances (Y-axis) across all graph models, along with training time (X-axis) and GNN parameter size (indicated by color). Note that there exist graph models yielding substantial improvements in both performance and efficiency compared to those generated by widely-used AR2N modeling [29].
B.2 Additional Results regarding Obs 2.
In this section, we provide additional experimental results supporting Section 4.2, which examines how modeling table rows can lead to markedly different outcomes depending on the specific downstream task. Figures 9 - 13 show the results on the remaining 10 tasks. The results for user-repeat and user-ignore are visualized in Figure 3 of the main paper.
Figure 9: Additional results on rel-f1 regarding Obs 2.
Figure 10: Additional results on rel-avito regarding Obs 2.
Figure 11: Additional results on rel-event regarding Obs 2.
Figure 12: Additional results on rel-stack regarding Obs 2.
Figure 13: Additional results on rel-trial regarding Obs 2.
B.3 Additional Results regarding Obs 3.
In this section, we present additional results to supplement Section 4.3. Figures 14 - 24 provide the case study results on the remaining 11 tasks. The case study result for user-attendance is provided in Figure 4 in the main paper.
Figure 14: Top performing graph models for driver-top3.
Figure 15: Top performing graph models for driver-dnf.
Figure 16: Top performing graph models for driver-position.
Figure 17: Top performing graph models for user-ignore. Note that the users_friends table has two FKs (user and friend) both referencing the users table.
Figure 18: Top performing graph models for user-repeat. Note that the users_friends table has two FKs (user and friend) both referencing the users table.
Figure 19: Top performing graph models for user-clicks.
Figure 20: Top performing graph models for user-visits.
Figure 21: Top performing graph models for ad-ctr.
Figure 22: Top performing graph models for user-ad-visit.
Figure 23: Top performing graph models for post-post-related. Note that the postLinks table has two FKs (RelatedpostId and postId) both referencing the posts table, and the posts table has three FKs (AcceptedAnswerId, OwnerUserId, and ParentId) all referencing the table itself.
Figure 24: Top performing graph models for study-outcome.
B.4 Additional Results regarding Obs 5.
In this section, we provide extra details and results to supplement Section 4.5. Given our computational constraints, we examine only the correlations for the top- and bottom-5% graph models—ranked by GraphSAGE (sum aggregation)—across three new predictors (i.e., graph neural networks): GraphSAGE (mean aggregation), GIN, and GPS.
As shown in Figure 25, while cross-GNN correlations are relatively low for the two tasks on the rel-f1 dataset (driver-top3 and driver-position) due to high variances in the performances, for the other tasks, cross-GNN correlations are high. The results for user-repeat, user-visits, and user-ad-visit are presented in Figure 6 in the main paper.
Figure 25: Additional results regarding Obs 5. Spearman correlations between different predictive GNNs—GraphSAGE with sum and mean aggregation (denoted as sum and mean), GIN, and GPS—are generally high (above 0.7) across the remaining tasks.
Appendix C Benchmark Detail of the RDB2G-Bench
C.1 Implementation Details of Baselines
In this section, we provide implementation details for the action-based search algorithms: Evolutionary Algorithm (EA), Bayesian Optimization (BO), and Reinforcement Learning (RL).
Evolutionary Algorithm. The Evolutionary Algorithm (EA) baseline employs a regularized evolutionary search strategy. The algorithm initializes a population by randomly sampling graph configurations and iteratively evolves this population through mutation and selection processes. At each iteration, a subset of individuals is selected through tournament selection, and mutations are applied using defined micro-actions to generate offspring. The oldest individual in the population is replaced by the best-performing offspring, maintaining diversity and continuous exploration.
Its key hyperparameter settings are summarized below:
  • Population Size: $\min(10,\text{budget})$
  • Tournament Size: $\min(10,\text{budget})$
  • Max Iterations: 1000
Bayesian Optimization. The Bayesian Optimization (BO) baseline employs an iterative search leveraging a surrogate Multi-Layer Perceptron (MLP) based on the BANANAS [37] implementation. The input embeddings represent graph structures converted into fixed-sized arrays based on selected edges.
The loss function follows the original BANANAS formulation:
$L=\text{mean}\left|\frac{y_{\text{pred}}-y_{\text{lb}}}{y_{\text{true}}-y_{\text{lb}}}-1\right|$ (8)
Initial embeddings are obtained through random sampling, and subsequent selections are guided by Expected Improvement, computed via Monte Carlo sampling with 50 samples. This iterative process continues until a defined budget limit or a maximum of 100 iterations is reached.
Its key hyperparameter settings are summarized below:
  • Surrogate Model: MLP (2 hidden layers, each with 32 units)
  • Dropout Rate: 0.1
  • Optimizer: Adam
  • Learning Rate: 0.001
  • Epochs per Iteration: 50
  • Batch Size: 32
  • Initial Sampling Size: $\min(10,\text{budget})$
  • EI MC Samples: 50
  • Max Iterations: 100
Reinforcement Learning. The Reinforcement Learning (RL) baseline utilizes a policy gradient approach with an RNN-based controller for the search process. The controller is implemented as a one-layer LSTM [15] with 32 hidden units. Input state embeddings are derived from the current graph structures encoded as fixed-size vectors.
Each training episode consists of up to five steps, during which the controller selects actions based on the policy generated from the RNN outputs. Rewards are computed based on performance improvements between sequential states, using a discount factor of 0.99. Training runs for up to 50 episodes or until a predefined evaluation budget is reached. Its key hyperparameter settings are summarized below:
  • Controller Model: LSTM (1 layer, 32 hidden units)
  • Optimizer: Adam
  • Learning Rate: 0.005
  • Max Steps per Episode: 5
  • Discount Factor ($\gamma$): 0.99
  • Episodes: 50
Appendix D Additional Results regarding Benchmark on the RDB2G-Bench Datasets
D.1 Performance Results on the RDB2G-Bench Datasets
In this section, we provide additional performance details of ten RDB-to-graph modeling methods, supplementing the summary presented in Table 1 and Figure 7(a) in the main paper. Details are presented in Tables 4 - 6 and Figure 26.
Table 4: Performance of ten RDB-to-graph modeling methods on the rel-event dataset under varying budget levels.
Task Name user-ignore (AUC-ROC (%) $\uparrow$) user-repeat (AUC-ROC (%) $\uparrow$) user-attendance (MSE $\downarrow$)
Methods Budget (%) Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 82.823 82.823 82.823 82.823 82.823 82.291 82.291 82.291 82.291 82.291 0.2375 0.2375 0.2375 0.2375 0.2375
Random 79.543 80.797 81.427 81.548 81.572 78.034 79.588 81.280 81.430 81.430 0.2446 0.2423 0.2396 0.2391 0.2389
AR2N 82.222 82.222 82.222 82.222 82.222 77.651 77.651 77.651 77.651 77.651 0.2445 0.2445 0.2445 0.2445 0.2445
GF 77.655 78.213 78.233 78.945 80.931 77.789 79.163 80.348 80.508 80.981 0.2447 0.2423 0.2399 0.2399 0.2385
GB 82.222 82.222 82.222 82.222 82.222 78.748 80.469 80.744 81.018 81.258 0.2432 0.2414 0.2409 0.2407 0.2400
GL 79.256 79.809 80.534 80.809 81.120 79.012 79.158 79.558 80.481 80.739 0.2423 0.2418 0.2415 0.2398 0.2391
EA 79.939 80.523 81.011 81.299 81.482 78.989 80.739 81.249 81.313 81.333 0.2425 0.2404 0.2395 0.2390 0.2389
BO 78.787 80.550 81.015 81.162 81.288 77.516 80.360 80.831 81.120 81.424 0.2449 0.2403 0.2396 0.2394 0.2392
RL 79.474 80.376 80.527 80.719 80.963 79.341 80.703 80.930 81.040 81.094 0.2474 0.2421 0.2412 0.2410 0.2409
LLM 80.541 80.541 81.146 81.751 81.751 78.841 79.537 80.857 80.941 80.941 0.2400 0.2397 0.2397 0.2397 0.2397
LLM-CoT 82.222 82.222 82.222 82.222 82.222 81.259 81.259 81.259 81.259 81.259 0.2394 0.2394 0.2387 0.2387 0.2387
Table 5: Performance of ten RDB-to-graph modeling methods on the rel-avito dataset under varying budget levels.
Task Name user-clicks (AUC-ROC (%) $\uparrow$) user-visits (AUC-ROC (%) $\uparrow$)
Methods Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 67.931 67.931 67.931 67.931 67.931 66.332 66.332 66.332 66.332 66.332
Random 66.594 66.856 67.098 67.151 67.302 64.444 65.169 65.521 65.708 65.877
AR2N 64.660 64.660 64.660 64.660 64.660 65.971 65.971 65.971 65.971 65.971
GF 66.978 67.096 67.096 67.096 67.096 66.318 66.318 66.318 66.318 66.318
GB 66.561 66.791 66.791 66.791 66.791 66.332 66.332 66.332 66.332 66.332
GL 65.770 66.183 66.468 66.468 66.483 64.700 65.090 65.118 65.124 65.124
EA 66.778 66.944 66.958 67.204 67.232 65.352 65.397 66.047 66.116 66.122
BO 66.388 67.069 67.170 67.442 67.584 64.823 65.004 65.380 65.924 66.191
RL 66.341 66.771 66.801 66.902 66.983 63.840 64.228 64.379 64.963 65.249
LLM 66.066 66.308 66.648 67.092 67.323 66.246 66.246 66.246 66.246 66.246
LLM-CoT 66.597 66.597 66.597 66.597 66.597 66.246 66.246 66.246 66.246 66.246
Task Name ad-ctr (MSE $\downarrow$) user-ad-visit (MAP (%) $\uparrow$)
Methods Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 0.0389 0.0389 0.0389 0.0389 0.0389 3.6816 3.6816 3.6816 3.6816 3.6816
Random 0.0399 0.0397 0.0396 0.0396 0.0396 3.2824 3.5017 3.6084 3.6468 3.6487
AR2N 0.0397 0.0397 0.0397 0.0397 0.0397 3.6610 3.6610 3.6610 3.6610 3.6610
GF 0.0402 0.0394 0.0394 0.0394 0.0394 3.6453 3.6453 3.6453 3.6453 3.6453
GB 0.0392 0.0392 0.0392 0.0392 0.0392 3.6610 3.6610 3.6610 3.6610 3.6610
GL 0.0412 0.0406 0.0406 0.0406 0.0406 2.2905 2.8499 3.0970 3.0989 3.1171
EA 0.0404 0.0400 0.0398 0.0395 0.0395 2.6709 3.3097 3.4353 3.4391 3.4432
BO 0.0399 0.0396 0.0394 0.0392 0.0392 3.0748 3.2248 3.4774 3.6241 3.6759
RL 0.0411 0.0406 0.0405 0.0403 0.0403 2.5915 2.8723 3.2453 3.2475 3.3514
LLM 0.0397 0.0397 0.0397 0.0397 0.0397 3.6610 3.6610 3.6610 3.6610 3.6610
LLM-CoT 0.0394 0.0394 0.0394 0.0394 0.0394 3.6610 3.6634 3.6655 3.6655 3.6679
Table 6: Performance of ten RDB-to-graph modeling methods on the rel-stack dataset (left) and the rel-trial dataset (right) under varying budget levels.
Task Name post-post-related (MAP (%) $\uparrow$) study-outcome (AUC-ROC (%) $\uparrow$)
Methods Budget (%) Budget (%)
1% 2% 3% 4% 5% 1% 2% 3% 4% 5%
Best 12.040 12.040 12.040 12.040 12.040 70.913 70.913 70.913 70.913 70.913
Random 11.751 11.803 11.835 11.867 11.900 70.178 70.338 70.372 70.375 70.407
AR2N 10.823 10.823 10.823 10.823 10.823 68.091 68.091 68.091 68.091 68.091
GF 11.903 11.903 11.903 11.903 11.903 69.572 69.572 69.572 69.572 69.572
GB 11.165 11.165 11.165 11.165 11.165 69.413 69.413 69.413 69.413 69.413
GL 10.555 10.555 10.555 10.555 10.555 69.490 69.490 69.490 69.490 69.490
EA 11.786 11.897 11.946 11.954 11.954 70.639 70.771 70.772 70.772 70.772
BO 11.860 11.888 11.888 11.888 11.888 70.376 70.376 70.376 70.376 70.376
RL 11.706 11.776 11.826 11.863 11.863 69.959 69.959 69.959 69.959 69.959
LLM 11.111 11.386 11.386 11.386 11.386 68.556 68.556 68.556 68.556 68.556
LLM-CoT 10.964 10.964 10.964 10.964 10.964 68.608 68.608 68.608 68.608 68.608
Figure 26: Performance details of ten RDB-to-graph modeling methods on each of the 12 predictive tasks. The performances are computed under varying budget levels, corresponding to the number of graph models evaluated.
D.2 Analysis of Evaluation Time
In this section, we present a detailed analysis of evaluation time costs for each RDB-to-graph modeling method, which complements the results summarized in Figure 7(b) in the main paper. Details are provided in Table 7.
Table 7: Analysis of evaluation costs for each method. Total time is computed as the sum of evaluation time and run time, encompassing on-the-fly evaluation and search method costs. All times are averaged over 10 runs.
Method Total Time (s) Run Time (s) Ratio (Total/Run)
Random 319016.09 0.59 540705.24
Greedy Forward Search (GF) 104246.09 0.39 267297.67
Greedy Backward Search (GB) 216032.31 0.51 423592.76
Greedy Local Search (GL) 284439.64 0.34 836587.18
Evolutionary Algorithm (EA) 696925.57 132.62 5255.06
Bayesian Optimization (BO) 595853.52 1244.17 478.92
Reinforcement Learning (RL) 753865.31 17.66 42687.73
LLM baseline (LLM) 50861.62 3588.22 14.17
LLM-CoT baseline (CoT) 28847.97 2924.82 9.86
D.3 Analysis of LLM Interpretability
Our analysis shows that interpretability can be enhanced when the LLM-CoT baseline provides explicit reasoning traces for each selected graph transformation action. During the action planning phase, we prompted the LLM to explain its rationale behind each operation, allowing us to capture structured insights about how it interprets temporal and structural dependencies.
As shown below, the LLM did not merely select from predefined actions but offered justifications referencing temporal order and causal consistency. These intuitive rationales demonstrate that the LLM’s reasoning process is consistent with the underlying graph structure, helping us understand why each transformation was made.
<reasoning>
1. Current State Analysis:
- The last action restored the FK-PK relationship between results and races, which improved performance slightly
- This suggests that temporal relationships between tables are important for DNF prediction
- The key tables for DNF prediction appear to be results, qualifying, races, and drivers
2. Examining Remaining Relationships:
- The FK relationship between standings and drivers was removed
- A drivers past performance and championship position could correlate with their risk-taking behavior or car reliability
3. Proposed Action:
- Restore the FK-PK relationship between standings and drivers
- Rationale:
* Driver standings contain cumulative performance metrics (points, position, wins)
* The temporal nature of standings data could help capture driver form and reliability trends
* Similar to how restoring race relationships improved performance, driver standings relationship might capture important patterns
4. Expected Impact:
- Should improve models ability to capture driver-specific patterns over time
- Aligns with the positive results seen from restoring other temporal relationships
Therefore, I recommend using add_fk_pk_edge to restore the relationship between standings and drivers.
</reasoning>
Appendix E Further Discussions
E.1 Limitations
Graph Design Space Constraints. As described in Section 3.1 of the main paper, we define the graph design space based on the choices in selecting foreign key relations and modeling rows as nodes or edges. However, additional potential design choices, such as creating dummy tables [35] and advanced feature engineering techniques, are currently excluded due to computational limitations. Expanding the graph design space can present a promising direction for future work.
GNN Dependence. Currently, RDB2G-Bench provides precomputed evaluations only for four GNNs: GraphSAGE (sum aggregation), GraphSAGE (mean aggregation), GIN, and GPS. Yet, our performance generalization analysis suggests that our findings and the benchmark’s utility extend beyond these specific GNNs.
Limited Task Coverage. While our original goal is to cover all tasks available in RelBench [29], we have excluded some tasks from rel-stack and rel-trial due to extremely large graph design spaces and high computational costs. Future work may extend our dataset by covering more tasks.
E.2 Broader Impact
Our analysis using RDB2G-Bench confirms the importance of strategic graph modeling for relational databases, highlighting its academic and practical significance.
From a research perspective, by providing a unified and well-structured set of benchmarks, RDB2G-Bench significantly reduces the evaluation cost for researchers and facilitates reproducible comparisons across methods. This enables more efficient and robust validation, thereby accelerating progress in RDB-to-graph modeling research.
Furthermore, leveraging RDB2G-Bench can significantly enhance RDB-to-graph modeling methods, enabling industries such as finance, healthcare, and e-commerce to improve efficiency and predictive performance on critical tasks. For example, the financial sector can enhance fraud detection accuracy, while the healthcare sector may develop more precise models for predicting patient outcomes, both directly benefiting from optimized graph models.
E.3 Analysis of GNN Depth
Following the default configuration of RelBench [29], we fixed the number of GNN layers to two in all experiments. This configuration provides a consistent and computationally efficient setup, and our ablation study further confirms that it is also empirically well-justified. As shown in Tables 8-9, the 2-layer GNN achieves the best average performance across the tasks while also exhibiting a strong correlation with the 3-layer GNN in terms of the top and bottom ranked graph configurations.
We adopt this configuration as it offers a consistent and efficient experimental setup. However, deeper GNNs could potentially better exploit longer relational paths, especially in databases with large-radius schemas. Exploring such depth variations remains a promising direction for future work.
Table 8: Average performance with different GNN depths (1–3 layers).
RDB Task Name (Metric) 1-layer 2-layer 3-layer
rel-f1 driver-dnf (AUC-ROC ↑) 71.05 ± 0.98 71.76 ± 1.39 70.87 ± 1.10
rel-avito user-clicks (AUC-ROC ↑) 64.03 ± 0.85 64.96 ± 1.03 64.88 ± 1.34
rel-event user-attendance (MAE ↓) 0.249 ± 0.011 0.249 ± 0.011 0.251 ± 0.010
Table 9: Correlation between 2-layer and 3-layer GNN configurations (top and bottom 10%).
RDB Task Name Pearson Spearman
rel-f1 driver-dnf 0.851 0.816
rel-avito user-clicks 0.852 0.805
rel-event user-attendance 0.963 0.883
Appendix F Prompt Design
In this section, we describe the prompts used for our LLM-based baseline implementation, introduced in Section 5.1.
Figure 27: Prompt template for main process.
Figure 28: Prompt template for action document.
Figure 29: Prompt template for error feedback.
Figure 30: Prompt template for performance feedback.
Figure 31: Prompt template for Chain-of-Thought reasoning.